OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 36?

OpenStudy (johnweldon1993):

\[\large f(x) = x^2 - 36\] \[\large f^{-1}(x) = ?\] We replace the 'x' with the 'y' and solve for 'y' again \[\large y = x^2 - 36\] \[\large x = y^2 - 36\] \[\large y^2 = x + 36\] \[\large y = \sqrt{x + 36}\] so \[\large f^{-1}(x) = \sqrt{x + 36}\]